OpenStudy (anonymous):

The container that holds the water for the football team is 1/3 full. After pouring out 6 gallons of water, it is 1/9 full. How many gallons can the container hold?

OpenStudy (texaschic101):

1/3x - 6 = 1/9x -6 = 1/9x - 1/3x -6 = 1/9x - 3/9x -6 = - 2/9x -6/(-2/9) = x -6 * -9/2 = x 54/2 = x 27 = x check.. 1/3(27) - 6 = 1/9(27) 9 - 6 = 3 3 = 3 (correct) tank holds 27 gallons